# OptiMUS-0.3

### v1

@@
Ali AhmadiTeshnizia , Wenzhi Gaob, Herman Brunborgb Shayan Talaeia, Madeleine Udella,b
@@
Optimization problems are pervasive in sectors from manufacturing and distribution to healthcare. However, most such problems are still solved heuristically by hand rather than optimally by state-of-the-art solvers because the expertise required to formulate and solve these problems limits the widespread adoption of optimization tools and techniques. We introduce a Large Language Model (LLM)-based system designed to formulate and solve (mixed integer) linear programming problems from their natural language descriptions. Our system is capable of developing mathematical models, writing and debugging solver code, evaluating the generated solutions, and improving efficiency and correctness of its model and code based on these evaluations. OptiMUS-0.3 utilizes a modular structure to process problems, allowing it to handle problems with long descriptions and complex data without long prompts. Experiments demonstrate that OptiMUS-0.3 outperforms existing state-of-the-art methods on easy datasets by more than 12% and on hard datasets (including a new dataset, NLP4LP, released with this paper that features long and complex problems) by more than 8%.
@@ 1 Introduction
Optimization problems are common in domains from operations and economics to engineering and computer science. Major advances in optimization algorithms over the last several decades have led to reliable and efficient optimization methods for a wide variety of structured optimization problems, including linear programming (LP) and mixed-integer linear programming (MILP), among many others. Unfortunately, optimization modeling — transforming a domain problem into a mathematical optimization problem — still requires expert knowledge. According to a recent survey, 81% of Gurobi’s commercial solver users have advanced degrees, with 49% of them holding a degree in operations research (Gurobi Optimization 2023). This expertise gap prevents many organizations from using optimization, even when it could significantly improve their operations. Examples include inventory management in supermarkets, patient operations in hospitals, transportation policies in small municipalities, energy management in local solar farms, and operations in small businesses or NGOs (Saghafian et al. 2015, Aastrup and Kotzab 2010, Yao et al. 2020, Shakoor et al. 2016). Automating optimization modeling would allow enterprises to improve efficiency using optimization techniques, even when they cannot afford to access optimization experts.
@@
  • Long problem descriptions. Realistic optimization problems can be exceedingly long and complex: for example, the documentation for the energy system problem described in Liu et al. (2023b) is 60 pages long. Unfortunately, LLMs have a limited context size, and even long-context models perform less well as the input context grows Liu et al. (2023b), Levy et al. (2024). Consequently, LLMs tend to make more mistakes as the length of the problem description increases and perform poorly on complex problems. How can we design an LLM-based optimization modeling system that keeps track of each individual constraint and term in the objective and ensures the model is complete and correct?
@@
This paper studies the feasibility of using LLMs to expand the reach of operations research techniques and analytic methods for decisionmaking by studying the performance of OptiMUS, an automated system we have developed that uses LLMs combined with traditional solvers to model and solve (mixed integer) linear programs (MILPs). OptiMUS advances the agenda of operations research when viewed as a meta-analytic tool to improve decisionmaking. We have chosen to begin with MILP modeling as MILP is one of the more classic, important, and widely used tools in the operations research toolbox. Moreover, examples of MILP problems, models, and solutions are plentiful and readily available. Our hope is that by making powerful solvers easily accessible to domain experts, OptiMUS can empower better, faster decisions that reduce costs and improve outcomes. Our studies of OptiMUS demonstrate both the promise and pitfalls of the LLM approach to quantitative decisionmaking. More broadly, through ablation studies, this paper clarifies the information flow needed to perform optimization modeling: in other words, it provides empirical evidence for what kind of education is effective to teach (LLMs) optimization modeling.
We call the version of OptiMUS studied in this paper OptiMUS-0.3 to distinguish it from previous versions that have been described in the arXiv draft AhmadiTeshnizi et al. (2023) and the conference publication Ahmaditeshnizi et al. (2024). Compared to these previous versions, OptiMUS-0.3 uses an LLM-based pipeline instead of the more flexible but error-prone agent-based system of OptiMUS-0.2. Our findings indicate that the increased flexibility in OptiMUS-0.2 led to more errors, outweighing the benefits at the current LLM capabilities. Additionally, nondeterministic errors made maintaining and improving the system challenging. Unlike OptiMUS-0.1, OptiMUS-0.3 uses a modular approach, using only relevant information for each constraint. OptiMUS-0.3 also incorporates Retrieval-Augmented Generation (RAG) and error correction, which were absent in both previous versions.
@@
  • We identify 18 important (non-exclusive) types of MILP problems, including scheduling, routing, production planning, and inventory management, which we use as labels for optimization problems. We curate a new dataset, NLP4LP, a comprehensive open-source dataset of 252 optimization problems that ensures adequate coverage of all these problem types, by combining problems from existing datasets with new problems. This dataset also includes problems that range in difficulty and description length, including real-world problems that are an order of magnitude longer that the problem instances in other MILP modeling datasets. Table 1 compares NLP4LP to existing datasets and Section 3 describes NLP4LP.
  • We develop a modular, LLM-based agent to model and solve optimization problems, which we call OptiMUS-0.3. OptiMUS-0.3 employs a connection graph that allows it to process each constraint and objective independently, which allows the system to solve problems with long descriptions and large data files without excessively long prompts.
  • We develop several new modules designed to improve the performance of OptiMUS-0.3 and study their impact, including:
@@
      Self-reflective error correction. We ask OptiMUS to assess its own confidence in its output, and fall back to a more powerful LLM or to user feedback when OptiMUS is unsure its answer is correct.
@@
      Retrieval-augmented generation. We search through a database of constraints and their formulations to find ones similar to the task at hand. We include a few of these as examples in the prompt which OptiMUS passes to the LLM.
    3. 3.
      Advanced optimization modeling techniques. We teach OptiMUS advanced optimization modeling techniques by prompting it to identify important structures in the problem, including structure of problems and solver features such as special ordered sets, and to use those structures to model and solve the problem more efficiently. We observe consistent improvement in terms of scalability and solution time after proper structure is identified.
  • We perform ablation studies of the components of our framework to demonstrate which elements improve performance and which can actually degrade performance. We find OptiMUS-0.3 beats the previous state-of-the-art methods on existing datasets by over 12% and on our more challenging test dataset by 8%.
@@ 2 Related Work
Optimization problems are mathematically defined by an objective function and a set of constraints. For example, an MILP can be written mathematically as
$\underset{\{x_{j}\}}{\text{minimize}}\quad$ $\sum_{j=1}^{n}c_{j}x_{j}$
subject to $\sum_{j=1}^{n}a_{ij}x_{j}{~{}(\leq,=,\geq)~{}b_{i}},i=1,\ldots m$
$l_{j}\leq x_{j}\leq u_{j},j=1,\ldots,n$
$x_{j}\in\mathbb{Z},j\in\mathcal{I}$
An optimization workflow consists of 1) formulating an optimization problem in mathematical form by identifying its objective and constraints, and then 2) solving the realization of problem from real data, generally using code that calls an optimization solver.
@@
Formulation is often a challenging task even for experts in optimization. Different formulations can lead to significantly different solving times and enable the use of different solvers or solution techniques (Boyd and Vandenberghe 2004). One important skill for an optimization expert is to identify assumptions or relaxations that allow for casting the problem as a well-studied problem type, such as MILP, which enables the use of well-developed optimization solvers. Crafting such an efficient formulation often requires specialized knowledge (Zohrizadeh et al. 2020, Low 2013, Roubíček 2020, Luo et al. 2010, Krarup and Pruzan 1983).
@@
Given the formulation, an optimization expert must choose a solver. Each solver has its own interface and functionalities (Achterberg 2019, Diamond and Boyd 2016, CPLEX User’s Manual 1987). However, the user manuals for these solvers are often hundreds of pages, making them difficult to fully understand and use.
@@
NLP4LP Hard (Ours) 912.3 $\pm$ 498.2 67 (18) $\checkmark$
Progress in LLMs.
Recent progress in Natural Language Processing (NLP) has led to the development of large language models (LLMs) useful for tasks such as answering questions, summarizing text, translating languages, and coding (OpenAI 2023, Touvron et al. 2023, Chowdhery et al. 2022, Wei et al. 2023, Gao et al. 2023a, Borgeaud et al. 2022). Connections to other software tools extend the reach and accuracy of LLMs, as demonstrated by plug-ins for code writing and execution (Paranjape et al. 2023, Wei et al. 2023). (Yang et al. 2023) use LLMs to directly generate solutions to optimization problems without calling traditional solvers through prompt optimization to improve performance. The approach is limited to small problems since the performance of LLMs degrades as the input context grows, even for explicitly long-context models (Liu et al. 2023a).
Chatbots for Optimization.
In a recent paper, (Chen et al. 2023) developed a chatbot to help users detect and fix infeasible optimization problems expressed in Pyomo code and servers as an AI assistant rather than as a solver. (Li et al. 2023) designed a chatbot to answer natural-language queries about an optimization model. (Alibaba Cloud 2022) also developed a chatbot to facilitate optimization modeling, but there is no public paper or documentation available on it.
Benchmark-driven Optimization Modeling.
More closely related to our approach, Ramamonjison et al (2023) introduced a dataset of 1101 natural language representations of LP problems. They proposed a two-stage mapping from the natural-language representation to the problem formulation using an intermediate representation. Ramamonjison et al (2022) designed a system to simplify and improve the modeling experience for operations research, but did not offer an end-to-end solution. Xiao et al. (2023) presented a multi-agent cooperative framework called Chain of Experts (CoE) to automatically model and program complex operation research (OR) problems, and evaluated it on NL4Opt and another more complex dataset, ComplexOR, introduced in that paper. Concurrent with the preparation of this paper, Tang et al. (2024) introduced a new approach to OR modeling with LLMs that uses fine-tuning on a semi-synthetic dataset rather than prompt optimization or agentic models, which they call ORLM (Operations Research Language Model). The numerical performance of their method is comparable to OptiMUS, but the work is complementary, as the fine-tuning methodology of ORLM could be combined with the agentic framework developed in the present work to improve performance further.
OptiMUS-0.3 is a variation of the previous optimization modeling systems we have built. OptiMUS-0.1 AhmadiTeshnizi et al. (2023) uses a set of unit tests generated by the LLM and revised by the user to evaluate and fix the generated code. If the code does not run or fails the tests it is passed to the LLMs along for the relevant error code for revision until it is fixed. OptiMUS-0.2 Ahmaditeshnizi et al. (2024) uses a structured sequence of LLM agents to effectively model and solve optimization problems. First, the natural language representation of the problem is preprocessed into a structured problem. Then, a team of agents iteratively augments the structured problem with a connection graph, mathematical formulations of each clause, and code for each clause, until the problem is solved. OptiMUS-0.3 uses the modular connection graph from OptiMUS-0.2, the pipeline from OptiMUS-0.1, and new error correction methods to achieve state-of-the-art performance.
Traditional MILP solvers generally benchmark against the MIPLIB benchmark, which offers a diverse collection of MILP problems in standard form. Unfortunately, most of these problems are not associated with a natural-language description, and so cannot be used to study optimization modeling as we do in this paper.
3 Dataset
Resource AllocationProduction PlanningSchedulingBlendingWorkforce ManagementSelectionSizingRoutingInventory ManagementPackagingPricingBalancingPortfolio OptimizationNetwork FlowCoveringCuttingFittingLayout0050505050100100100100150150150150200200200200250250250250300300300300350350350350324324324324179179179179808080805555555537373737363636363333333328282828171717171515151513131313131313131010101088884444333333332222ValuesOperational Areas
Figure 1: NLP4LP covers all common optimization problem types
Manufacturing and ProductionSupply Chain ManagementFood and BeverageTransportation and LogisticsHealthcare and MedicalRetail and E-commerceEnvironmental and SustainabilityAgriculture and ForestryScience and ResearchEnergy and Power SystemsFinance and BankingSports and EntertainmentGovernment and Public SectorEducationHuman ResourcesTelecommunicationsMarketing and MediaAerospace and Defense002020202040404040606060608080808010010010010012012012012014014014014014214214214210210210210278787878727272726868686853535353303030302929292919191919151515151515151515151515131313131212121212121212333333332222ValuesIndustry Sectors
Figure 2: NLP4LP covers most common optimization problem domains
23456780050505050100100100100150150150150ValuesRatings Count
Figure 3: Most problems have around 4 labels.
We introduce NLP4LP, a comprehensive open-source dataset of 344 optimization problems. Our goal in creating NLP4LP is to provide the community with examples they can use to design optimization modeling tools using natural language systems and to assess their quality. It consists of two sub-datasets: 1) An easy dataset which contains only LP problems with short descriptions and scalar parameters, and 2) a hard dataset which contains both LP and MILP problems with longer descriptions and multi-dimensional parametrs (Table 1).
Given the complexity of the task, and the components of each instance (description, sample feasible data, and solution code), gathering high-quality optimization instances is expensive. Unlike traditional ML datasets, NLP4LP does not provide a training set with the same distribution as the test set. As the cost of collecting each instance is high, we have prioritized the task of benchmarking by emphasizing the test set rather than reducing the size of the test set to provide training data. Consequently, NLP4LP is partitioned into 1) a development (dev) set of 22 problems (12 easy and 10 hard) and 2) a test set of 332 problems (277 easy and 55 hard). We designed the partition intending that the dev set be used to develop optimization modeling tools, while the test set be used to evaluate such tools. The dataset is available publicly at https://nlp4lp.vercel.app/
Optimization modeling is used in many domains. To ensure that NLP4LP comprehends a wide variety of use cases, we first create a list of important types of optimization problems, such as scheduling, cutting, routing, blending, and packing, and a list of common application domains such as sports, government, retail, agriculture, and energy. We tag each problem NLP4LP with all relevant labels, and continue to gather instances until the dataset contains at least two instances per label. We collect the instances from textbooks, lecture notes, research papers, and existing datasets Bertsimas and Tsitsiklis (1997b), Williams (2013), Nace (2020), Xiao et al. (2023), Ramamonjison et al (2023) Figure 3 shows that most problems have around four labels, and some have as few as two or as many as eight. Figures 1 and 2 provide histograms showing the number of problem instances for each type and domain; resource allocation problems are the most common.
Real-world problems are often much longer and more complicated than textbook problems. For the test set, we have included ten long real-world problems from scientific papers. These problems are an order of magnitude longer than the instances in the other datasets (in terms of number of variables and constraints).

### v2

@@
Ali AhmadiTeshnizia , Wenzhi Gaob, Herman Brunborgb Shayan Talaeia, Connor Lawlessa, Madeleine Udella,b
@@
Optimization problems are pervasive in sectors from manufacturing and distribution to healthcare. However, most such problems are still solved heuristically by hand rather than optimally by state-of-the-art solvers because the expertise required to formulate and solve these problems limits the widespread adoption of optimization tools and techniques. We introduce a Large Language Model (LLM)-based system designed to formulate and solve (mixed integer) linear programming problems from their natural language descriptions. Our system is capable of developing mathematical models, writing and debugging solver code, evaluating the generated solutions, and improving efficiency and correctness of its model and code based on these evaluations. OptiMUS-0.3 utilizes a modular structure to process problems, allowing it to handle problems with long descriptions and complex data without long prompts. Experiments demonstrate that OptiMUS-0.3 outperforms existing state-of-the-art methods on easy datasets by more than 22% and on hard datasets (including a new dataset, NLP4LP, released with this paper that features long and complex problems) by more than 24%.
@@ 1 Introduction
Optimization problems are ubiquitous in domains from operations and economics to engineering and computer science. While major advances in optimization algorithms over the last several decades have led to reliable and efficient optimization methods for a wide variety of structured optimization problems, including linear programming (LP) and mixed-integer linear programming (MILP), optimization modeling — transforming a domain problem into a mathematical optimization problem — still requires expert knowledge. According to a recent survey, 81% of Gurobi’s commercial solver users have advanced degrees, with 49% of them holding a degree in operations research (Gurobi Optimization 2023). This expertise gap prevents many organizations from using optimization, even when it could significantly improve their operations. Examples include inventory management in supermarkets, patient operations in hospitals, transportation policies in small municipalities, energy management in local solar farms, and operations in small businesses or NGOs (Saghafian et al. 2015, Aastrup and Kotzab 2010, Yao et al. 2020, Shakoor et al. 2016). Automating optimization modeling would allow enterprises to improve efficiency using optimization techniques, even when they cannot afford to access optimization experts.
In this paper, we take a first step towards an automated system that can democratize access to advanced optimization modeling expertise. While the ultimate aim of this line of research is to enable real-world stakeholders without any optimization background to formulate and solve their own complex decision-making problems, we study a necessary precursor to such a system - an automated system to accelerate the workflow of existing optimization practitioners. We specifically focus on helping novice and intermediate optimization users develop, tune, and implement (mixed-integer) linear optimization problems, a flexible and broad class of optimization problems that have been applied successfully in a wide range of real-world use cases such as scheduling and production planning. Our work aims to improve the productivity of optimization experts, akin to the productivity gains realized by software engineers with systems such as GitHub CoPilot, and lay the foundation for more robust systems that can ultimately cater to non-expert users. Given that our focus is on existing optimization practitioners, we assume for the remainder of this paper that the user has already identified a problem for which MILP is an effective tool and can describe the problem in a level of detail such that modeling is possible. Aiding users to find the correct modeling paradigm (e.g., constraint programming, conic optimization) and supporting them in concretizing vague problems remains an important open problem for future research.
@@
  • Long problem descriptions. Realistic optimization problems can be exceedingly long and complex: for example, the documentation for the energy system problem described in Holzer et al. (2023) is 60 pages long. Unfortunately, LLMs have a limited context size, and even long-context models perform less well as the input context grows (Liu et al. 2023, Levy et al. 2024). Consequently, LLMs tend to make more mistakes as the length of the problem description increases and perform poorly on complex problems.
@@
This paper studies the feasibility of using LLMs to expand the reach of operations research techniques and analytic methods for decision making by studying the performance of OptiMUS, an automated system we have developed that uses LLMs combined with traditional solvers to model and solve (mixed integer) linear programs (MILPs). Our studies of OptiMUS demonstrate both the promise and pitfalls of the LLM approach to quantitative decision making, and help clarify the effectiveness of LLMs for optimization modeling as of 2024. More broadly, through ablation studies, this paper shows how structured reasoning pipelines provide a clear pathway to address current limitations and enhance the reliability of LLM-based systems in optimization.
@@
  • We curate a new dataset, NLP4LP, a comprehensive open-source dataset of 355 optimization problems that ensures adequate coverage of a wide variety of problem types, by combining problems from existing datasets with new problems. Our dataset includes problems that range in difficulty and description length, including real-world problems that are an order of magnitude longer than the problem instances in other MILP modeling datasets. Table 1 compares NLP4LP to existing datasets and Section 3 describes NLP4LP. Our dataset also includes many problems that are less likely to have been used to train existing LLMs, reducing the risk of data leakage compared to previous optimization modeling datasets.
  • We develop a modular, LLM-based agent to model and solve optimization problems, which we call OptiMUS (or, occasionally, OptiMUS-0.3 to distinguish it from prior versions). OptiMUS-0.3 employs a connection graph that allows it to process each constraint and objective independently, which allows the system to solve problems with long descriptions and large data files without excessively long prompts.
  • We develop several modules designed to improve the performance of OptiMUS and study their impact, including:
@@
      Self-reflective error correction. We ask OptiMUS to evaluate and correct its output, and assess its own confidence in its output, allowing it to fall back to a more powerful LLM or to user feedback when OptiMUS is unsure its answer is correct.
@@
      Advanced optimization modeling techniques. We teach OptiMUS advanced optimization modeling techniques by prompting it to identify important structures in the problem, including structure of problems and solver features such as special ordered sets, and to use those structures to model and solve the problem more efficiently.
  • We perform ablation studies of the components of our framework to demonstrate which elements improve performance and which can actually degrade performance. We find OptiMUS-0.3 beats the previous state-of-the-art methods on existing datasets by over 22% and on our more challenging test dataset by 24%, and GPT-4o alone by over 38% and 40% on easy and challenging datasets respectively.
  • We have developed a publicly available web application that allows practitioners to start using OptiMUS. This webapp also illustrates an initial approach to human computer interaction for LLM-powered optimization modeling. Notably, the interface makes it easy for the human to audit and correct the LLM at each stage of modeling, and allows the LLM to communicate uncertainty to the user in order to solicit corrections when needed.
An initial version of this work was published in a conference proceeding (Anonymous 2024) that introduced a smaller initial version of the NLP4LP dataset, and a less robust LLM agent-based system for modeling optimization problems that we will refer to as OptiMUS-0.2. The new enlarged version of the NLP4LP dataset which accompanies this paper has been open-sourced and includes a modular design that enables future researchers to improve on different sub-problems within our modular optimization framework such as parameter extraction, modeling, or coding. Compared to the earlier version, the modular design of OptiMUS-0.3 substantially improves the robustness of the system for more complicated modeling problems. New modules that enable the system to leverage advanced solver functionality, such as special-order-set constraints, can dramatically speed up the solution time for the associated optimization problems. Additional LLM techniques in OptiMUS-0.3 such as self-reflective error correction improve performance. Finally, this paper is the first to present the OptiMUS webapp.
@@ 2 Related Work
Formulating an optimization model is often a challenging task even for experts in optimization. Different formulations can lead to significantly different solving times and enable the use of different solvers or solution techniques (Boyd and Vandenberghe 2004). One important skill for an optimization expert is to identify assumptions or relaxations that allow for casting the problem as a well-studied problem type, such as MILP, which enables the use of well-developed optimization solvers. Crafting such an efficient formulation often requires specialized knowledge (Zohrizadeh et al. 2020, Low 2013, Roubíček 2020, Luo et al. 2010, Krarup and Pruzan 1983).
@@
Given the formulation, an optimization expert must choose a solver. Each solver has a distinct interface and capabilities, with associated benefits and downsides (Achterberg 2019, Diamond and Boyd 2016, CPLEX User’s Manual 1987). However, the user manuals for these solvers are often hundreds of pages, making them difficult to understand and use.
Progress in LLMs.
Recent progress in Natural Language Processing (NLP) has led to the development of large language models (LLMs) useful for tasks such as answering questions, summarizing text, translating languages, and coding (OpenAI 2023, Touvron et al. 2023, Chowdhery et al. 2022, Wei et al. 2023, Gao et al. 2023a, Borgeaud et al. 2022). Connections to other software tools extend the reach and accuracy of LLMs, as demonstrated by plug-ins for code writing and execution (Paranjape et al. 2023, Wei et al. 2023). Yang et al. (2023) use LLMs to directly generate solutions to optimization problems without calling traditional solvers through prompt optimization to improve performance. The approach is limited to small problems since the performance of LLMs degrades as the input context grows, even for explicitly long-context models (Liu et al. 2023).
Chatbots for Optimization.
In a recent paper, Chen et al. (2023) developed a chatbot to help users detect and fix infeasible optimization problems expressed in Pyomo code and servers as an AI assistant rather than as a solver. Li et al. (2023) designed a chatbot to answer natural-language queries about a supply chain optimization model. Alibaba Cloud (2022) also developed a chatbot to facilitate optimization modeling, but there is no public paper or documentation available on it. Lawless et al. (2024) explored the use of LLMs to allow users to customize a simple constraint programming model in the context of meeting scheduling, but does not model more general MILP optimization problems.
Benchmark-driven Optimization Modeling.
More closely related to our approach, Ramamonjison et al (2023) introduced a dataset of 1101 natural language representations of LP problems. They proposed a two-stage mapping from the natural-language representation to the problem formulation using an intermediate representation. Ramamonjison et al (2022) designed a system to simplify and improve the modeling experience for operations research, but did not offer an end-to-end solution. Xiao et al. (2023) presented a multi-agent cooperative framework called Chain of Experts (CoE) to automatically model and program complex operation research (OR) problems, and evaluated it on NL4Opt and another more complex dataset, ComplexOR, introduced in that paper. Astorga et al. (2024) present a Monte-Carlo Tree Search (MTCS) based approach that decomposes the modeling problem into stages (e.g., modeling variables, then objectives, then constraints, etc.) and uses MTCS to explore the space of plausible models. As mentioned in the paper, this approach is not directly comparable to OptiMUS as it outputs a set of functionally distinct models as opposed to a single formulation and answer. Concurrent with the preparation of this paper, Tang et al. (2024) introduced a new approach to OR modeling with LLMs that uses fine-tuning on a semi-synthetic dataset rather than prompt optimization or agentic models, which they call ORLM (Operations Research Language Model). Yang et al. (2024) also explore a synthetic data generation and fine-tuning approach for a broader class of optimization problems. These fine-tuning approaches are complementary to our work, as any progress towards more capable LLMs for optimization modeling can be combined with our modular prompting framework.
Traditional MILP solvers generally benchmark against the MIPLIB benchmark (Gleixner et al. 2021), which offers a diverse collection of MILP problems in standard form. Unfortunately, most of these problems are not associated with a natural-language description, and so cannot be used to study optimization modeling as we do in this paper.
3 Dataset
As part of this paper, we have curated a comprehensive open-source dataset of 355 optimization problems we call NLP4LP. Our goal in creating NLP4LP is to provide the community with examples they can use to design optimization modeling tools using natural language systems and to assess their quality. The problems in NLP4LP are partitioned into 1) an easy dataset that contains only LP problems with short descriptions and scalar parameters, and 2) a hard dataset that contains both LP and MILP problems with longer descriptions and multi-dimensional parameters (Table 1). Unlike other modeling datasets, each instance in the NLP4LP dataset is accompanied by associated code to run the instance, and ground-truth intermediary representations of the problem (i.e., extracted parameters and targets, a list of clauses of the problem represented in natural language, , and code, and a solution and optimal value for given problem data). This additional structure enables future researchers to investigate and improve different modules of the OptiMUS framework such as parameter extraction or code implementation, and even test these improvements with the OptiMUS webapp, without implementing the full software stack from scratch. Compared to the initial version of the NLP4LP dataset presented in (Anonymous 2024) this version contains more instances from a wider breadth of domains and more complex problems from published papers, and includes additional information for each instance, such as code, as discussed earlier. The dataset is available on Hugging Face and in the supplementary material.
Given the complexity of the task, and the components of each instance (description, sample feasible data, and solution code), gathering high-quality optimization instances is expensive. Unlike traditional ML datasets, NLP4LP does not provide a training set with the same distribution as the test set. As the cost of collecting each instance is high, we have prioritized the task of benchmarking by emphasizing the test set rather than reducing the size of the test set to provide training data. Consequently, NLP4LP is partitioned into 1) a development (dev) set of 22 problems (12 easy and 10 hard) and 2) a test set of 332 problems (277 easy and 55 hard). We designed the dev/test partition intending that the dev set be used to develop optimization modeling tools, while the test set be used to evaluate such tools. In our experiments and system design, we did not use any part of the test set for development, ensuring an unbiased evaluation of the models.
Optimization modeling is used in many domains. To ensure that NLP4LP comprehends a wide variety of use cases, we first created a list of important types of optimization problems, such as scheduling, cutting, routing, blending, and packing, and a list of common application domains such as sports, government, retail, agriculture, and energy. We tagged each problem NLP4LP with all relevant labels, and continued to gather instances until the dataset contains at least two instances per label. Instances in our dataset are drawn from several sources: some were created by our research team, inspired by problems in textbooks, lecture notes, and research papers, and some are drawn from existing datasets on LP modeling (Bertsimas and Tsitsiklis 1997, Williams 2013, Nace 2020, Xiao et al. 2023, Ramamonjison et al 2023). Most problems have around four labels, and some have as few as two or as many as eight. For more information on problem types and domains refer to Appendix 24. Real-world problems are often much longer and more complicated than textbook problems. For the test set, we have included ten long real-world problems from scientific papers. These problems are an order of magnitude longer than the instances in the other datasets (in terms of number of variables and constraints) and seriously challenge the modeling capabilities of existing LLM systems (including ours). Table 1 shows that our dataset contains instances that are substantially longer than the problem instances in other MILP modeling datasets.
One major concern in evaluation of LLM-based systems is data leakage: has the LLM been trained on the data in the purported test set? This issue is serious and has been known to contaminate the results of several important studies (Oren et al. 2023). In our study, we take several measures to mitigate the risk of leakage, balancing this concern against the goal of broad coverage in our test set.
  • Our dataset is guarded by Captcha, and only authenticated Hugging Face users can access it after agreeing to terms and conditions. These safeguards prevent the dataset from being used to train future LLMs.
  • Many problems in the dataset are inspired by textbook examples and recent publications, yet their text is entirely original. Moreover, the sources of these questions do not include solutions (either in or code). We verified that the content of these questions does not appear anywhere on the internet using a plagiarism detection tool, which yielded a mean originality score of $84.9\%$ and a median originality score of $100\%$. See Appendix 24 for details.
  • Some LPs in our dataset are drawn from other datasets created prior to the training cutoff for many models (Ramamonjison et al 2023). However, these instances are rephrased and modified in our dataset. Moreover, reporting results on these problems is standard practice (Meta AI 2024, Google 2024). None of the MILPs, which represent the most challenging problems in our dataset, are sourced from existing datasets.
@@ 3 Dataset
NLP4LP Hard (Ours) 912.3 $\pm$ 498.2 68 (18) $\checkmark$